Grade 11IIT JEE Entrance Exam

the lengths of the perpendiculars from the focus and the extremities of a focal chord of a parabola on the tangent at the vertex form
  1. an AP
  2. a GP
  3. an HP
  4. none of these

Arnav Jaitly
11 Years agoGrade 11

1 Answer

Nishant Vora
11 Years ago
Answer is 3 HP

Let one coordinate of focal chord is P(at2, 2at)
So ohter cordinate will be Q(a/t2,-2a/t)
Also S(a,0) is the focus

Now PS=PM=l1 and QS=QM=l2
Now l1= distance from directrix = a + at2
Now l1= distance from directrix = a + a/t2

Now find out \frac{2*l1*l2}{l1+l2}
this will come equal to ‘a‘(semi-major axis)

#its an important result you can remember also.